OpenStudy (anonymous):

The height of a rock at time t which is thrown vertically on the height 46 feet is given by the formula s(t) = −t 2 + 20t+ 44. What is the maximum height of the rock? When does it hit the ground? What is the impact speed?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

im just stuck on when the rock his the ground and the impact i figured out the height

OpenStudy (compassionate):

@jfernandes Testing...

OpenStudy (photon336):

Not sure I'm wondering if we can just take the derivative of S(t) and set it equal to zero. \[\frac{ dS }{ dt } = -2t+20 \] \[\frac{ dS }{ dt }_{0} \] \[-2t+20 = 0 \] \[\frac{ -20 }{ -2 } = 10 = t \] I feel that when the velocity is zero the will reach it's maximum height. so we plug this back into the original equation. \[s(2) = -(2)^{2}+20(2)+44 = 80ft\] When the rock hit's the ground the position will be zero so \[s(t) = -t ^{2}+20t+44 = 0 \] i'm thinking we can apply this formula \[-b \pm (b ^{2}-4ac)^{.5}/2a\]

OpenStudy (anonymous):

so far the maximum height is 144 and i checked to see if it was right and it was

OpenStudy (photon336):

oh wait.. i plugged in the wrong number

OpenStudy (anonymous):

so i the derivative at 0 t = 10 then i plugged it back in and got -(10)^2 + 20(10) + 44 = 144 which is the max height

OpenStudy (anonymous):

so now im wondering how to i find te value of t when the rock hits the ground

OpenStudy (photon336):

\[-(10)^{2} + 20(10)+44\] -100+200 = 100+44 = 144 @reeses Yeah I plugged in 2 instead of 10

OpenStudy (photon336):

when the rock hits the ground @reeses the height will be zero, vertical height

OpenStudy (danjs):

the acceleration is constant, the flight is symmetrical, so time to max height, is same as time down

OpenStudy (photon336):

-b +/- (b^2-4ac)^0.5/2a ((20)^2-4(-1)(44))^0.5 -20-19.6 = -40/2(-1) = 20 seconds

OpenStudy (danjs):

is that 46 the initial height, i didnt see that?

OpenStudy (photon336):

i'm thinking that to find the velocity we would plug t back into the equation for velocity. dS/dt = -2t+20 t = 20 -2(20)+20 = -20 m/s i'm guessing the negative sign means it's downward @DanJS I did not see that, oh... yeah .. the initial height was 46ft but @reeses said the answer for height was correct I'm guessing?

OpenStudy (anonymous):

sweet! i got it now so it was -(t-22)(t+2)=0 t = 22 t = -2 which -2 isnt right so t = 22 is right and its the right answer

OpenStudy (anonymous):

now how about the impact

OpenStudy (danjs):

Yeah, in the prob, the position function at time 0 is s(t) = 44, but it says starts at 46,

OpenStudy (anonymous):

i know its throwing me off

OpenStudy (photon336):

@reeses if you found the time at impact , I think you can plug it back into the equation, the derivative dS/dt which gives velocity

OpenStudy (danjs):

yes that is for when it hits the ground, it has to fall that extra 44 ft on the way down, with a initial velocity the same as the initial velocity it was thrown upwards

OpenStudy (anonymous):

well the time of impact to the ground is 22

OpenStudy (danjs):

yeah so 10 to top , 10 back to the h=44, and must be 2 more sec down

OpenStudy (danjs):

if you use the initial velocity and the accel and height, should get that time

OpenStudy (photon336):

how would you set it up?

OpenStudy (anonymous):

so i looked up the answer and its 4 feet/sec i have no idea how thats the answer

OpenStudy (danjs):

s(t) = -t^2 + 20t + 44 [ft] v(t) = -2t + 20 [ft/s] a(t) = -2 [ft/s^2] --------------------------------

OpenStudy (danjs):

The initial height s(0) = 44 ft, thrown up with an initial velocity of v(0) = 20 ft/s undergoes a constant acceleration of -2 ft/s^2

OpenStudy (danjs):

kimematic equation to get time from 20 ft/s to 0 ft/s with constant accel of -2 is 10 seconds

OpenStudy (anonymous):

but how did the text get 4 feet/sec

OpenStudy (danjs):

s(10) = max height time to fall from that max height to h=0 must be 22 s

OpenStudy (danjs):

sorry 12s

OpenStudy (anonymous):

could it be a typo?

OpenStudy (danjs):

the ground inpact speed will be greater (neg direction) than the initial velocity because it is accelerating downwards longer than deceleration to zero speed,

OpenStudy (danjs):